A normal distribution has a mean of μ = 80 with σ = 12. Find the following probabilities. (a) p(X > 83) (b) p(X < 74) (c) p(X < 92) (d) p(71 < X < 89)
Answer #1

a)

Since μ 80 and σ = 12 we have X-μ P(X>83 ) = P (X-1 > 83-80 ) = P (-> - 83-80)-83-80 Since Z-z-μ and 83-80-0.25 we have: P(X>83 ) = P ( Z > 0.25) Step 3: Use the standard normal table to conclude that: P (Z > 0.25) 0.4013

b)

c)

d)